OpenStudy (anonymous):

Input the equation of the given line in slope-intercept form. The line through (1, 0) and parallel to x - y = 7.

terenzreignz (terenzreignz):

If it's parallel to x - y = 7, it must have the same slope with x - y = 7, right?

OpenStudy (jackellyn):

To find the slope of x-y =7 put it in y =mx + b form, where m is the slope.

terenzreignz (terenzreignz):

Not too cooperative, are we? :) x - y = 7 can be rewritten as y = x - 7 The slope, therefore, is...? (come on, work with us here :) )

OpenStudy (jackellyn):

haha, yeah we are trying to teach you! :)

OpenStudy (anonymous):

hahah i got it now thanks!